OpenStudy (anonymous):

Find the point in the first quadrant on the graph of y=x^2 that is closest to the point (0,4)

OpenStudy (perl):

use d( (x,y) , (0,4) ) = sqrt ( (x-0)^2 + (y-4)^2) and substitute y = x^2

OpenStudy (perl):

distance between a point on graph y=x^2 and (0,4) is d = sqrt ( x^2 + (x^2 - 4)^2 )

OpenStudy (perl):

d = sqrt ( x^2 + x^4 - 8x^2 + 16 )

OpenStudy (perl):

d = sqrt ( x^4 -7x^2 + 16)

OpenStudy (perl):

so a trick is actually to square both sides first d^2 = x^4 - 7x^2 + 16

OpenStudy (anonymous):

Okay, got that. Do I need to plug in the points now?

OpenStudy (perl):

you need to solve d'(x)

OpenStudy (perl):

2*d *d'(x) = 4x^3 -14x d'(x) = (4x^3 -14x)/ (2d)

OpenStudy (perl):

set that equal to zero (4x^3 -14x) / (2d) = 0 4x^3 - 14x = 0 the denominator =0 is also a critical pt, but the denominator is strictly greater than zero

OpenStudy (perl):

2x(2x^2 -7) = 0 x^2 = 7/2 x = + - sqrt (7/2)

OpenStudy (anonymous):

And how could I solve for y?

OpenStudy (perl):

we know y = x^2

OpenStudy (perl):

( sqrt (7/2), 7/2)

OpenStudy (anonymous):

So square x right?

OpenStudy (anonymous):

Ha! Okay got it! Thanks!!

OpenStudy (perl):

did that make sense

OpenStudy (anonymous):

Yes a lot actually! You factored the numerator and equaled it to zero to solve for x, and then plugged it into the given equation.

OpenStudy (perl):

right, and we don't care about the denominator 2d , because it is always greater than zero

OpenStudy (perl):

technically we skipped the part which shows this is the minimum . you show that d'(x) > 0 for 0< x < sqrt(7/2) and d'(x) < 0 for x < sqrt(7/2)